OpenStudy (anonymous):

What is the limit as x approaches infinity, (sqrt(x^2+9x+1)-x)?

OpenStudy (anonymous):

Now if the same rule applies I should end up with (sqrt(x^2(1+9/x+1/x^2))-x)

OpenStudy (jamesj):

Multiply this expression by ( sqrt(x^2+9x+1) + x ) / ( sqrt(x^2+9x+1) + x ) And you'll see that this simplifies. By the way, what is the answer intuitively of this limit?

OpenStudy (anonymous):

i get (10x +1)/2x?

OpenStudy (jamesj):

No. (sqrt(x^2+9x+1)-x) . ( sqrt(x^2+9x+1) + x ) / ( sqrt(x^2+9x+1) + x ) = sqrt( x^2+9x+1)^2 - x^2 ) / ( sqrt(x^2+9x+1) + x ) , because the num was a diff of squares = ( x^2 + 9x + 1 - x^2 ) / ( sqrt(x^2+9x+1) + x ) = ( 9x + 1 ) / ( sqrt(x^2+9x+1) + x ) Now ... how do you handle this expression?

OpenStudy (anonymous):

yea, i forgot that x cancels out so it would leave 9x+1/2x...therefore giving me 9/2...i believe

OpenStudy (jamesj):

Yes, correct.

OpenStudy (jamesj):

'good answer' if you don't mind.

OpenStudy (valpey):

Is this the question? \[\lim_{x\rightarrow\infty}\sqrt{(x^2+9x+1)} -x\]

OpenStudy (zarkon):

yes

OpenStudy (valpey):

\[\infty\]

OpenStudy (zarkon):

no...it is 9/2

OpenStudy (jamesj):

No Valpey. It's 9/2.

OpenStudy (jamesj):

Follow the calculations above and you'll see why.

OpenStudy (valpey):

\[\lim_{x\rightarrow\infty}\sqrt{(x^2+9x+1)} - x\] \[\lim_{x\rightarrow\infty}\frac{(\sqrt{(x^2+9x+1)} - x)(\sqrt{(x^2+9x+1)} + x )}{(\sqrt{(x^2+9x+1)} + x )}\] \[\lim_{x\rightarrow\infty}\frac{(x^2+9x+1) - x^2}{(\sqrt{(x^2+9x+1)} + x )}\] \[\lim_{x\rightarrow\infty}\frac{9x+1}{\sqrt{(x^2+9x+1)} + x}\] \[\lim_{x\rightarrow\infty}\frac{9x+1}{x\sqrt{(1+9/x+1/x^2)} + x}\] \[\lim_{x\rightarrow\infty}(\frac{9x}{x\sqrt{(1+9/x+1/x^2)} + x}+\frac{1}{x\sqrt{(1+9/x+1/x^2)} + x})\] \[\lim_{x\rightarrow\infty}(\frac{9}{\sqrt{(1+9/x+1/x^2)} + 1}+\frac{1}{x\sqrt{(1+9/x+1/x^2)} + x})\] \[(\frac{9}{\sqrt{(1)} + 1}+0) = \frac{9}{2}=4\frac{1}{2}\]